OpenStudy (anonymous):

a pool for swimming laps is designed to be seven times as long as it is wide. Length = 7x Width = x If the areas of the sides and the bottom are 90m^2, what are the dimensions of the pool if the volume of water it can hold is to be maximised?

OpenStudy (ikram002p):



OpenStudy (ikram002p):

then area=7x^2+2xh+14xh=90 7x^2+(2x+14x)h=90 h=(90-7x^2)/16x volume(x)=7x^2*(90-7x^2)/16x (maximize this function)

OpenStudy (ikram002p):

\(v(x)=630x-49x^3\\ v'(x)=630-147x^2\\ 0=630-147x^2\\ x^2=(30/7)\\ x=\pm \sqrt (30/7)\) you could make a test and know that the positive one is the max :) i think its like this , hope dint make a typo or something :)

OpenStudy (anonymous):

Thank you. I tried before I checked answer and got the same after about 3 hours thought. I guess I am right. Thank you again